### v1

# Model Compression in Practice: Lessons Learned from Practitioners Creating On-device Machine Learning Experiences

Abstract.
On-device machine learning (ML) promises to improve the privacy, responsiveness, and proliferation of new, intelligent user experiences by moving ML computation onto everyday personal devices. However, today’s large ML models must be drastically compressed to run efficiently on-device, a hurtle that requires deep, yet currently niche expertise. To engage the broader human-centered ML community in on-device ML experiences, we present the results from an interview study with 30 experts at Apple that specialize in producing efficient models. We compile tacit knowledge that experts have developed through practical experience with model compression across different hardware platforms. Our findings offer pragmatic considerations missing from prior work, covering the design process, trade-offs, and technical strategies that go into creating efficient models. Finally, we distill design recommendations for tooling to help ease the difficulty of this work and bring on-device ML into to more widespread practice.
Efficient machine learning, machine learning model compression, on-device machine learning, interview study, interactive systems, design directions
$c$s: Human-centered computing Interactive systems and toolsccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligence
1. Introduction
Most modern machine learning (ML) models in production today occupy cloud servers of exceedingly more computational capacity and power than the device in your pocket. Yet for delivering intelligent user experiences, there is good reason to shift ML models onto our everyday personal computing devices.
On-device ML is the practice of storing, training, and running ML models on an individual’s device, such as a smartphone, tablet, or wearable. However, as today’s state-of-the-art models grow larger and larger in size (e.g., into the billions of parameters) (Villalobos et al., 2022; Giattino et al., 2022; Zhao et al., 2023; Stanford, 2023), efficiency remains the biggest barrier to on-device ML (Apple, 2023a; Developers, 2022). An arbitrary ML model placed on a mobile device can easily consume every available resource of the device, whether it be compute, memory, or battery. Creating efficient, on-device models brings new challenges to the ML development process.
In this paper we argue that research in efficient machine learning over the past decade has matured to the degree to where practitioners have an (albeit rapidly evolving) set of techniques to make on-device ML a reality. The key idea is to shrink, optimize, and compress models, while simultaneously maintaining their accuracy. To achieve this, practitioners develop strategies for how to best apply model compression techniques to minimize the amount of computational resources needed. Product designers, tool-builders, and ML practitioners today should be considering the benefits of on-device ML for their users:
First, on-device models can be an enormous win for personal privacy. While the industry standard practice today is to send a user’s (encrypted) data over a network to a commercially-owned server for ML inference, on-device ML cuts out the need for a server such that a user’s personal data never leaves their device. Beyond inference, pre-trained models can be fine-tuned on-device to adapt and personalize to an individual user’s preferences, while keeping those preferences private and local (Konečnỳ et al., 2016; Hu et al., 2020; Apple, 2023b).
Second, models on-device enable intelligent user experiences where they would not be possible otherwise. For example, computational photography models within mobile cameras run inference at high frame rates (e.g., 30 frames per second), which would not be possible if relying on the availability of a distant server.
Third, by going offline, on-device ML can create not only faster but also more portable experiences. Since no data is sent over a network, users can still interact with ML-powered features in situations without internet access or cellular service. This has the potential to broaden access of AI/ML features in more geographic regions, including rural and remote areas (Tomašev et al., 2020). By forgoing a network entirely, ML-powered features can run faster and remain responsive by alleviating network latency in an “offline” mode (Developers, 2022).
Lastly, removing the reliance on servers for ML inference has economic and environmental impact. On-device ML cuts out the cost of server, which may help individuals, non-profits, or smaller tech firms that previously could not afford server upkeep to now deliver ML-powered features to their users. By reducing society’s reliance on external servers, on-device ML may help reduce the high carbon footprint of the cloud (Monserrate, 2022).
In this paper, we seek to help the broader community adopt on-device ML by filling a crucial gap: while efficient ML research and algorithmic advances have progressed (Zhao et al., 2022; Gu et al., 2021; Zamzam et al., 2019; Dhar et al., 2021; Sehgal and Kehtarnavaz, 2019), there currently exists little pragmatic guidance (NVIDIA, 2023) in literature, online, in books, or otherwise, for newcomers wanting to create on-device ML users experiences. To curate pragmatic guidance, we interviewed 30 industry ML practitioners (research scientists, ML engineers, and product managers) who are uniquely experienced with designing, developing, and deploying on-device ML at scale. We capture the hard-won knowledge, insights, and tricks-of-the-trade that are traditionally omitted from technical ML publications to illustrate the current state-of-the-art. We also profile how industry practitioners adapt cutting-edge ML research into intelligent experiences running on peoples’ everyday devices. We contribute:
  • Tacit knowledge that 30 expert practitioners have developed around the design process, trade-offs, and strategies that go into deploying efficient on-device ML. Many decisions around optimization and compression strategy stem directly from user experience and product design.
  • A characterization of the key challenges practitioners face when creating efficient models. Example challenges include the tension between optimizing performance against model accuracy, and the necessity to work with hardware-level details.
  • Distilled design recommendations for interactive interfaces, tooling, and systems that show promise to help practitioners optimize models and ultimately proliferate on-device ML experiences.
We conclude by discussing where the broader human-centered AI/ML research, design, and practitioner communities can engage in efficient, on-device ML. We hope the results from this deep dive into a nascent area of ML user experience design helps spotlight its interdisciplinary importance and inspires others to contribute.
2. Related Work in Human-Centered ML and Model Optimization
Over the past decade, interview studies from human-computer interaction (HCI) have played an important role in capturing how AI/ML plays out in real-world practice (Amershi et al., 2019a, b; Holstein et al., 2019; Sambasivan et al., 2021; Liao et al., 2020; Bhatt et al., 2020; Yang et al., 2018). Interviews with practitioners are uniquely suited to capture the kinds of process-oriented insight, stories, and tricks-of-the-trade that are traditionally omitted from technical ML publications. Interviews also crucially give the broader public access to learn from AI/ML work that otherwise happens behind closed doors. In this paper, we aim to share new ideas for designing powerful on-device ML experiences, based on extensive product experience that is currently held by few. To situate our work in the broader landscape of HCI + ML research, we cover other interviews studies within diverse disciplines of AI/ML research, including user experience design, best practices for development, and interactive systems and tools.
2.1. User Experience Design for AI/ML
Notably, much of what the research community knows today about designing effective user experiences (UX) for AI/ML has been taught through the hard-won experience of real designers. For example, Yang et al. interviewed 13 UX practitioners with 4+ years of experience working on AI/ML products, to capture designers’ strategies for working with AI/ML (Yang et al., 2018). Zdanowska et al. interviewed both product teams and UX designers to explore how design collaboration happens during AI/ML product development (Zdanowska and Taylor, 2022). Windl et al. interviewed another 20 interaction designers to learn how designers adapt their their innovation process for AI/ML experiences (Windl et al., 2022). Again with experienced UX + AI/ML designers, Yildirim et al. conducted workshops around AI/ML innovation and found that designers are more impactful when allowed to design entire services—rather than fitting AI/ML into specific user-interfaces for existing products (Yildirim et al., 2022). Meanwhile, leading technology companies have distilled design wisdom from their own product teams into AI/ML design guidelines as public, educational resources (Wright et al., 2020). Examples include Apple’s Human Interface Guidelines on Machine Learning (gui, 2019a), Microsoft’s Human-AI Guidelines (Amershi et al., 2019b), Google’s People + AI Guidebook (gui, 2019b), and IBM’s Design for AI resources (des, 2019).
This body of work emphasizes that designing with machine learning requires new approaches from designers (Yang, 2018; Amershi et al., 2019b; Windl et al., 2022; Yildirim et al., 2022; gui, 2019b; des, 2019). Our paper contributes to this conversation by illustrating how power and performance issues of moving ML models on-device create real, tangible UX constraints for designers (for example, see Table 2)—and offering concrete design strategies for creating effective user experiences around those constraints (for example, see Section 5.4).
2.2. Profiling Best Practices for Building AI/ML
Regarding the engineering of AI/ML development, practitioner stories inform best-practices and new research agendas. Amershi et al. interview and survey AI/ML practitioners at Microsoft to gather best practices for production ML development and outline core differences between ML development and conventional software engineering (Amershi et al., 2019a). Researchers have studied how AI/ML teams collaborate across diverse technical roles (Passi and Jackson, 2018; Zhang et al., 2020; Piorkowski et al., 2021), or adopt new technology like AutoML (Wang et al., 2019, 2021b, 2021a). Sambasivan et al. used interviews to profile organizational struggles between balancing modeling work with data quality work (Sambasivan et al., 2021). Holstein et al. found key mismatches between academic AI/ML fairness concepts and the kinds of real fairness issues product teams grapple with (Holstein et al., 2019). Many other aspects of AI fairness, accountability, transparency, and ethics (FATE) have been found to be fundamentally sociotechnical issues where design decisions and process have enormous impact (Hong et al., 2020; Madaio et al., 2020; Modarres et al., 2018; Delgado et al., 2021; Madaio et al., 2022; Robertson et al., 2021; Rakova et al., 2021; Hopkins and Booth, 2021).
From a practical perspective within the ML community, existing work highlights the hidden technical debt within ML systems that incur massive amounts of engineering maintenance (Sculley et al., 2015). Some these ideas can be traced back before the proliferation of ML (Hand, 1998). More recently, related to our work is a recent technical survey on efficient methods for natural language processing (NLP) (Treviso et al., 2022) that aims to inspire new technical methods for efficiency in the age of ever-scaling language models; however, its primary focus is to only organize existing techniques specific to NLP applications. In our work, we profile on-device ML holistically, including its impact on product design, organization-level coordination, and best-practices for ML teams.
2.3. Interactive Systems for AI/ML Development
Extensive prior work in HCI, visualization, and ML has focused on interactive systems and tools to help ML practitioners overcome key challenges in their work. Examples include tools for: neural network architecture visualization and debugging (Wongsuphasawat et al., 2017), visualization for large ML dataset exploration (Google, 2021b, a) and data iteration (Hohman et al., 2020), model evaluation for classification tasks (Görtler et al., 2022; Amershi et al., 2015; Ren et al., 2016), visual analytics for deep learning interpretability (Hohman et al., 2018), fairness evaluation for finding disproportionately underperforming subpopulations of data (Bird et al., 2020; Cabrera et al., 2019), reliable error analysis and evaluation for NLP models (Wu et al., 2019; Ribeiro et al., 2020; Ribeiro and Lundberg, 2022), interactive visualization for counterfactual generation (Wu et al., 2021; Wexler et al., 2019), and interaction techniques to enhance ML and data science environments (Patel et al., 2008), such as computational notebooks (Kery et al., 2020) and other composable, shareable interfaces (Bäuerle et al., 2022). Today, on-device and efficient machine learning has little tooling available outside of code libraries to run compression algorithms (see applications within popular ML toolkits in Section 2.6). Our results from interviewing 30 experts provide strong designs recommendations for interactive tooling opportunities that would make on-device ML more achievable for everyone.
2.4. On-device Inference versus On-device Training
It is worthwhile to note a distinction between on-device inference and on-device training. On-device ML inference refers to running an ML model on any device with limited compute resources, such as limited compute, memory, or power. These models are pretrained, where learning happened offline on a different device or server. On-device ML training refers to either training a model from scratch, or fine-tuning a model for a few iterations, on a user’s device.
While work in on-device ML encompasses both paradigms, in our paper we focus on the far more common use case of model inference (deploying pretrained models on a user’s device), and leave on-device training for future work. Although on-device inference and training share many of the same core challenges around efficiency, training on-device models is generally considered harder, since training usually requires far more resources (Zhou et al., 2019). Thus, we do not attempt to cover the added complexities of on-device training in this work. For a survey on the current research and challenges around on-device learning instead, see: (Dhar et al., 2021; Lim et al., 2020; Zhou et al., 2019; Murshed et al., 2021).
2.5. Edge Computing
In recent literature, the challenges of on-device ML are often discussed in the context of edge computing (Zhou et al., 2019; Murshed et al., 2021; Chen and Ran, 2019). Edge computing typically refers to small embedded hardware, wearables, or internet-of-things (IoT) devices where at least some computation is performed on-device (the “edges” of a network) rather than by a central server (Murshed et al., 2021). In our research, we interview ML practitioners deploying to a variety of consumer devices, including edge hardware. We do not distinguish between device type except for where practices slightly differ for edge hardware, as in Section 5.4.4.
2.6. Existing Compression Resources
Beyond academic surveys detailing different compression techniques (Deng et al., 2020; Choudhary et al., 2020; Cheng et al., 2018), most practical guidance for model optimization is found within online tutorials and documentation for popular ML libraries. Examples include TensorFlow’s model optimization toolkit and blog post on quantization-aware training method (TensorFlow, 2020, 2018); PyTorch’s experimental support for quantization (PyTorch, 2018), sparsity (PyTorch, 2019), and it’s accompanying examples (PyTorch, 2023); Google’s quantization extension to Keras called QKeras (Google, 2019); Microsoft’s Neural Network Intelligence package and tool (Microsoft, 2021); Intel’s Neural Compressor library (Intel, 2020); and Apple’s DNIKit (Welsh, 2023). Other examples that target specific hardware include speeding up inference on FPGAs (Fahim et al., 2021) and compressing Core ML models to run on Apple platforms (Apple, 2023a). In terms of other community efforts, the appropriately named TinyML community has sprung up around this topic, which has published a book (Warden and Situnayake, 2019) and hosts community events and meetups for like-minded practitioners.
2.7. Compressing Large Language Models and Foundation Models
Within the past couple of years, model compression has garnered intense interest due to the public proliferation of large language models (LLMs) and foundation models (Bommasani et al., 2021). LLMs have moved beyond the research community and captured public interest and concern (Solaiman et al., 2023; Movva et al., 2023), as they have demonstrate new ways to interact with computers via natural language. The catch is that these models, by definition, are the largest in the world (e.g., billions of parameters (Giattino et al., 2022; Villalobos et al., 2022)). For all the reasons already mentioned, there is immense interest and activity of moving LLMs on-device—and progress is moving fast. The number of empirical works for compressing LLMs from this year alone would be difficult to track. However, we note that the same overall approaches for efficient machine learning that we discuss in this work (detailed in Section 3) are the same techniques being applied to LLMs. For interested readers in learning about the latest results for compressing LLMs, we refer to the follow surveys: (Treviso et al., 2023; Zhou et al., 2023; Xu and McAuley, 2023).
3. A Primer on Model Compression Techniques
To familiarize readers with model optimization in general, here we give a primer on common ML compression techniques. This background will help ground the study results and provide context for the remainder of the paper.
Model compression is a class of techniques used in on-device ML that optimizes varying components of a model to minimize their required computational resources. While it is not important to know every detail of every technique, it is useful to understand of the variety of techniques, how they work at a high-level, and how they can be combined together. In fact, some of the biggest savings can come from effectively combining multiple techniques (Han et al., 2016). This overview contains a brief description of each technique, and includes illustrations to build visual intuition. We only attempt to cover techniques mentioned by practitioners in our study. Each technique below is truly a family of techniques, each with many nuanced variations. For an in-depth review of the technical descriptions of compression techniques, see the following surveys: (Deng et al., 2020; Choudhary et al., 2020; Cheng et al., 2018).
3.1. Quantization
Convert the inputs, outputs, weights, and/or activations of a model from high-precision representations (e.g., fp32) to lower-precision representations (e.g., fp16, int32, int16, int8, and even int2). At a high-level, this coarsens a model. For a detailed survey of quantization-specific techniques and their variations, see  (Gholami et al., 2021).
3.2. Palettization (Weight Clustering)
Map the weights of a model to a discrete set of precomputed (or learned) values (Cho et al., 2022; Wu et al., 2018). Inspired by an artist’s palette, the idea is to map many similar values to one average or approximate value, then use those new values for computing inference. In this way, palettization is similar in spirit to algorithmic memoization, or more simply, a dictionary or a look-up table. Palettization can make a model smaller but does not make a model faster since it incurs look-up time.
3.3. Pruning (Sparsity or Network Sparsity)
Remove the least-important parts of a neural network to make it smaller (Hoefler et al., 2021). The motivation here is that modern ML networks are much more dense than is actually needed. Since networks can contain millions of parameters, removing some or many parameters may not impact the final accuracy. Pruning is a large family of techniques and strategies. For a detailed overview see (Hoefler et al., 2021).
3.3.1. Unstructured Pruning
A model may have neurons or weights that do not highly affect a model’s decision (Gholami et al., 2021). In unstructured pruning, the least important neurons or weights can be aggressively removed without affecting model accuracy (Hoefler et al., 2021). Unstructured pruning has been shown to shrink a model by 10–100x its original size (Hoefler et al., 2021). The major downside is that this leads to sparse matrix operations, which is a type of math that is slow on most hardware. So although pruned models are small, they may be nearly as slow as the full-sized model (Gholami et al., 2021).
3.3.2. Structured Pruning
Similar to unstructured pruning, the least important neurons or weights are identified, however, instead of removing (or zero-ing out) individual values, structured pruning removes entire structural elements, like channels or filters. Since structured pruning removes much bigger chunks of a model, less pruning will be possible without serious degradation in model accuracy (Gholami et al., 2021). However, by respecting the structure of the model, these pruned models maintain dense matrix operations, which make them fast on most hardware (Hoefler et al., 2021).
3.4. Distillation
Train a smaller model to mimic a larger model. Given a large (teacher) model that is already highly accurate at a task, the idea is to transfer the “knowledge,” or learned features, to a smaller (student) model by having the smaller model replicate the larger model’s output. Extending this further, given a newly trained student model, one can then apply other compression techniques to the student model for further optimization (Gholami et al., 2021; Polino et al., 2018). For a survey of distillation techniques see (Gou et al., 2021).
3.5. Efficient Neural Architectures
Some model architectures are specifically designed to be small and efficient. One of the most prominent examples in recent years is MobileNets (Howard et al., 2017; Sandler et al., 2018), with others like ShuffleNets (Zhang et al., 2018), or EfficientNets (Tan and Le, 2019). All of these examples are designed to improve the efficiency of convolutions, which are vital to fitting computer vision models on-device (Choudhary et al., 2020). Efficient architectures can be compressed with techniques like quantization or pruning, to create an even smaller model (Choudhary et al., 2020).
3.6. Dynamic Models
When thinking about ML applications, most people think of a single model with fixed inputs and outputs. Dynamic models differ in that they take into account the the differing prediction difficulty of each data instance. In other words, based on the individual data point to run inference on, a model may adapt its preprocessing steps, computational path, or model choice all together (Zhu et al., 2021). While many types of dynamic models exist, we only discuss a selection of techniques that are mentioned later on.
3.6.1. Early Exit Models
Allow for completing a prediction without the need to pass through the full model. Motivated by the fact that some data points are easier to predict than others, early exit models run intermediate feature representations through additional output layers to check if the prediction is sufficiently correct. For example, one way to determine whether or not to early exit or continue through a network is to check its prediction confidence: if a model is already confident about the current prediction, early exit, and if not, continue to the next prediction layer. At a high-level, these are models with checks throughout an architecture, such that inference can complete at any check before the final output.
3.6.2. Gated Models
Train a smaller, approximate model whose prediction decides whether or not to invoke a larger model. Similar to early exit models, some data points are easier to predict than others. A fast, smaller model gates a larger model, for example by again using the confidence of the initial prediction. At a high-level, these are systems of models that decide whether or not to spend extra compute if a prediction is still unknown.
4. Interview Study Methodology
4.1. Study Protocol
To capture emerging practices around efficient machine learning, we conducted semi-structured interviews (Boyce and Neale, 2006; Knott et al., 2022) with 30 ML practitioners at Apple to study how they approach model compression in their own projects. Our interview questions are outlined in Appendix A. We gave participants ample time to flexibility speak to their specific work and express other topics beyond our question set that they felt were important to doing good efficient machine learning work (Knott et al., 2022). The interviews took place between March and July 2022 with each conversation lasting from 45 minutes to 1 hour. For all interviews, one researcher lead the interview questions, while another took detailed notes. Where a participant approved, we also recorded conversations to refer back to during analysis. No compensation was given, since all participants were salaried employees. At the end of the study, we briefed participants and their teams on our results. Our study protocol was approved by an internal IRB.
4.2. Participants and Recruitment
As discussed in Section 2, in-depth interview studies with ML practitioners are uniquely suited to capture experts’ tacit knowledge for the purpose of generating new publicly-available guidance (Amershi et al., 2019a, b; Holstein et al., 2019; Sambasivan et al., 2021; Liao et al., 2020; Bhatt et al., 2020; Yang et al., 2018). Our organization is one of the few places in the world that contains a concentration of efficient machine learning experts, so we first reached out to known experts. We then used a snowball sampling strategy to reach a broader network of people involved in efficient machine learning at Apple, including other compression experts, ML practitioners who own a production model but are not compression experts themselves, and tooling engineers who build frameworks for model compression. To balance different perspectives, we sought practitioners working on ML for different domains (e.g., vision, text, and sensor data) and different use cases. As the study evolved, we also chose new participants to help fill-in our biggest areas of uncertainty. Our participants, listed in Section 5.1, include ML research scientists, engineers, and managers spanning a wide breadth of application domains. A natural saturation occurred when participants began to repeat know-how we had already recorded and began to only suggest participants we had already included.
4.3. Qualitative Data Analysis
The first two authors conducted an iterative thematic analysis method to group common motivations, challenges, and best practices of model compression into categories (Gibbs, 2007). As we conducted more interviews, we continually discussed and updated these categories to reflect our new findings. Each participant’s data and transcripts were independently reviewed and manually coded by the first two authors at least twice using inductive coding (Thomas, 2003). In total, we spent 30+ hours interviewing participants and collected 23,500+ words of additional notes outside of the audio transcripts. The final categories were split into two main sections: Section 5 “Study Results” and Section 6 “Design Opportunities for Efficient Machine Learning Interface.” The composition of these two sections was formed by ordering major categories and hierarchically grouping within categories when relevant.
4.4. Methodological Limitations
While we found that learning from expert practitioners was tremendously valuable, any interview study has limits for generalizability. Interview studies suffer from smaller population samples than other methods, for example, a survey. We chose to hold in-depth exploratory conversations with each participant—which a survey cannot support as well (Knott et al., 2022). On the other hand, we did not directly observe participants conducting their efficient machine learning work, which limits the specificity of our findings. It was not possible to observe participant’s direct work or model artifacts due to the sensitivity of their work. Another concern is bias from a participant’s role in the domain (Boyce and Neale, 2006), as well as the power dynamics between interviewer and interviewees (Kvale, 2006). This study was conducted solely within one organization, therefore practitioners may hold organization-specific beliefs and practices (Schein, 1990). Despite our conscious efforts to recruit across ML application domains, we noticed a skew towards vision-based applications on images, video, and 3D scene data. This makes sense from an industry perspective, as computer vision applications typically rely on exceptionally large ML models. Lastly, these interviews were conducted prior to the public rise of LLMs in late 2022 (Movva et al., 2023). As discussed in subsection 2.7, the same overall approaches for efficient machine learning are currently being applied to LLMs and other foundation model modalities.
Taking these limitations together, we caution readers to not consider the advice we detail from participants universally generalizable. However it is with confidence that we share the rich pragmatic advice, guidance, and experience these 30 experts have to offer.
5. Study Results
5.1. Participants and Emergent Personas
We summarize our participants in Section 5.1. Of the 30 people we interviewed, participants had an average of 7.1 (max 12) years experience with ML. Participants had 4.1 (max 10) years on average experience with efficient ML. Our participants had diverse breadth of expertise across domains and mapped roughly to four roles: 6 ML Managers, 4 Research Scientists, 13 ML Engineers, and 7 Software Engineers. For the purpose of understanding practitioners roles relative to efficient machine learning, we found three distinct personas emerged that best describe our participants:
Table 1. A summary of the 30 participants from the interview study. Participants are grouped by three emergent personas: Compression Experts (E), ML Practitioners (P), and Toolkit Engineers (T). Participants indicated the number of years they have spent working on machine learning (light blue squares), and of those years how many have been spent working on model compression and optimization specifically (dark blue squares).
A full-page table divided into three sections, each describing one of the personas of the study participants. Each row represents a participant, and lists their years of experience, job title, and what ML application they focus on.
ID Experience (Compression / ML Years) Job Title (Manager ✓) ML Application
Compression Expert
[2pt/2pt] E1 10/12 ML Manager (✓) Efficient ML research
E2 7/12 ML Manager (✓) 3D computer vision
E3 5/12 ML Manager (✓) Efficient ML research
E4 5/10 ML Engineer (✓) 3D computer vision
E5 9/9 Research Scientist Efficient ML research
E6 6/8 ML Engineer (✓) Efficient computer vision
E7 5/8 ML Engineer (✓) 3D computer vision
E8 2/8 ML Manager (✓) Efficient ML research
E9 5/6 ML Engineer (✓) Efficient computer vision
E10 5/6 Research Scientist Efficient ML research
E11 5/5 ML Engineer Efficient ML research
E12 5/5 Research Scientist 3D computer vision
E13 3/5 ML Engineer 3D computer vision
ML Practitioner
[2pt/2pt] P1 6/10 ML Manager (✓) Efficient computer vision
P2 4/10 ML Engineer Machine translation
P3 3/10 Research Scientist (✓) ML sensing
P4 4/9 ML Engineer Machine translation
P5 5/8 Software Engineer Computer vision
P6 1/6 ML Engineer Multi-modal ML
P7 1/6 ML Engineer Multi-modal ML
P8 2/5 ML Engineer ML fairness
P9 3/4 ML Engineer ML hardware
P10 3/4 ML Engineer 3D computer vision
P11 1/4 Software Engineer (✓) Multi-modal ML
Tooling Engineer
[2pt/2pt] T1 5/10 ML Manager (✓) Efficient ML tooling
T2 3/6 Software Engineer Efficient ML tooling
T3 2/6 Software Engineer ML fairness
T4 4/4 Software Engineer Efficient ML algorithms
T5 3/3 Software Engineer Efficient ML algorithms
T6 2/2 Software Engineer Efficient ML tooling
  • Compression Experts (E1–E13 in Section 5.1): Seasoned research scientists and experienced engineers that lead on-device ML initiatives, develop new compression techniques, and consult on machine learning optimization efforts. Given the amount of experience these people have, they often manage or lead teams. Example: An ML research scientist who has a PhD in model compression and is developing novel techniques for model optimization.
  • Machine Learning Practitioners (P1–P11 in Section 5.1): ML engineers and data scientists that build and deploy models on-device where certain optimization budgets must be met. These people use compression as a means to an end rather than solely studying efficient machine learning. Example: An ML engineer needs to get a model’s size under 1MB while maintaining high accuracy.
  • Tooling Engineers (T1–T6 in Section 5.1): Engineers and developers that focus on building frameworks, infrastructure, and tools for efficient machine learning. Example: A software engineer that builds and maintains an organization-wide tool to help others compute efficiency metrics.
Throughout the paper, we use representative quotes from participants to illustrate the main findings from the study.
5.2. The State of Efficient Machine Learning
Creating an efficient machine learning model is a “large design and constrained optimization problem” [E5], where practitioners have to weigh decisions between many model performance metrics, such as model storage size, inference latency, power usage, device temperature, and model behavioral metrics such accuracy, precision, recall, and others. To many readers, this characterization may sound like a job for an automated optimization algorithm. However, a crucial finding of our results is that automation is not yet possible, do to the degree of both ML and product design human expertise that goes into creating on-device applications. Experts emphasized there is no single solution, “silver bullet” [E12], “turnkey solution” [E5], or “golden recipe” [E7] for successfully building efficient machine learning models. Partially, this is due to the rapidly moving-target of new state-of-the-art model architectures and new ML applications that may require a custom approach.
Table 2. Key aspects of on-device ML efficiency that impact user experience.
A full-width table with two columns: user perceivable metrics on the left and negative user impact on the right. The user perceivable metrics column contains colored icons to indicated metrics like model storage size as a box, power usage as an empty battery, and device temperature as a flame.
User Perceivable Metrics Negative User Impact
[Uncaptioned image] Model storage size Modern deep learning models can easily grow to gigabytes. Devices have finite storage, so a model must not occupy so much space that it interferes with a user storing their own content.
[Uncaptioned image] Power usage Users expect their devices to have long battery lives, but a resource-intensive model can quickly drain a battery. Addressing battery drain is particularly challenging for older devices with lower battery capacity and health.
[Uncaptioned image] Device temperature ML models that consume lots of computational power at once can heat up a user’s device to where it is uncomfortably warm to the touch. Heat can also trigger thermal throttling, where the device slows down to avoid over-heating.
[Uncaptioned image] Inference latency Taking a long time to run inference can be frustrating for a user waiting on the result, and can make the experience feel unresponsive.
[Uncaptioned image] FPS (frames per second) A special case of inference latency. When processing live inputs (e.g., video), this latency can create delays that makes the output appear choppy or unresponsive.
[2pt/2pt] All metrics: model tested during high device resource load Resources are shared. A user’s device may be running many things simultaneously, including multiple ML models. A single model that demands too many resources will slow everything down.
“[Compression] works super well sometimes, but sometimes totally fails. Each project is too dependent, recipes tend to only work in some situations.” — P2, P6, P7
The expertise for creating efficient ML models is currently held by a relatively small subset of ML practitioners, where “knowledge is handed-down from the few who know how to do it successfully,” [E3]. These people are referred to as artisans and “based on earlier experience, know it’s possible” [P10] to create efficient models. Reducing a model’s size by half, or by an order of magnitude, takes additional time and considerations [E7]. “It’s like black magic or the dark arts, but you will get better with time,” [E7]. Readers may recognize these “dark arts” or “artisan” language from earlier days of contemporary ML (Patel et al., 2008; Hill et al., 2016) before any real widespread knowledge base was established. We hope these study results support growing such a knowledge-base for efficient machine learning, and it’s role within a holistic ML process:
[Efficient machine learning] means a lot of things to different people. Specific techniques on models weights help reduce size, but to get an efficient model comes from more careful design of the loss function, the system, which parts should and should not be modeled with ML.” — E5
In practice, building efficient ML models describes the process of either building a new model from scratch or modifying an existing model to shrink it enough to fit performance constraints, such as model size, latency, power consumption, and even heat [P3]. Newcomers often gravitate towards the latest model compression techniques from literature as a first-step, while experts who consult on these projects encourage taking a step back to consider the entire data and model lifecycle: “Don’t do [model compression] blindly. Don’t do [model compression] in a rush,” [T1]. “Philosophically, [you] need to look at whole problem, not as an afterthought,” [E3].
In addition to barriers of overall experience, efficient ML work differs perhaps most significantly in how much it requires a deep understanding of hardware details. Low-level hardware details are not something that typical ML or software engineering usually needs to consider. “[ML engineers] feel a bit intimated, and people don’t feel like they understand this stuff well,” [T2].
To this end, we think there is rich opportunity to help share and inform practice. Our interview study results characterize how practitioners design on-device machine learning experiences (5.3), decide model budgets (5.4), invest in compression (5.5), and evaluate compressed models to balance accuracy versus performance and avoid compression artifacts (5.6). We conclude with distilled design recommendations for tooling that could address the challenges practitioners face and bring efficient machine learning practices mainstream (Section 6). Throughout the results we highlight summarized takeaway strategies in call-out boxes.
5.3. The Metrics of Efficient, On-Device Models
For many practitioners we spoke to, efficient machine learning is of immediate importance to their work in developing new product features. Efficiency is a deciding point on which features get approved to ship to users, so efficient ML enables user experiences that otherwise simply could not exist [T5, P3, E12, P2, E9].
“Who cares if we can detect X if your model takes too much space? It will never make it to engineering requirements.” — P3
As an umbrella term, efficiency encompasses many metrics. Common metrics that practitioners linked to user experience are summarized in Table 2. Some may be familiar with metrics commonly considered in ML workflows, such as latency and model storage size. Other metrics may come as more foreign to an ML engineer accustomed to deploying their model on a far-off server, e.g., device temperature. When a user is holding and carrying around a device all day, metrics like heat and battery life impact are crucial.
The prioritization of efficiency metrics depends on the model type and its intended usage. While different practitioners we spoke to were focused on optimizing different aspects of efficiency, they all had a shared concept of a model budget.
5.4. Deciding Model Budget
A model budget encompasses thresholds for speed, accuracy, size, or the amount of any specific resource a model is allowed to consume. Model budgets are created individually per model, and are based as much in product and UX design as they are in device constraints.
5.4.1. Budgeting by Technical Feasibility
For new and novel ML applications, it is simply unknown on the onset how computational intensive a model might be. Practitioners refer to reported metrics from related ML research and an organization’s other models to sketch out an initial budget:
“At the beginning of the design, [we do a] ‘back of envelope calculation’ of where things need to be as honest as possible to what’s realistic. At the beginning, you’re more focused on accuracy. Over time there is refinement.” — E7
A common refinement practice is outlined in Figure 1. First, ML engineers will work on a “feasibility” model simply to see if the ML task “is possible at all” [P3]. Once the feasibility model works at sufficient accuracy (Figure 1A), engineers work on an architecture search to find a smaller, more efficient model that achieves the accuracy goal (Figure 1B) [P3, E7]. Next, the budget is refined based on the potential to shrink the model using compression: “given [an] accuracy goal, what is the biggest percentage reduction possible?” [E11]. Some teams do this reduction estimation on their own, while others bring in ML compression experts [T5]. Since the compressed-model budget (Figure 1C) is often estimated before engineers embark on model compression, the budget remains open to refinement, subject to product design constraints. Ultimately if an ML-powered feature requires a strict model budget, the model will not ship until engineers have found a way to reach those thresholds [P3].
Figure 1. Practitioners start with a feasibility model (A), which is a model of any size that demonstrates that the ML task works. Next, an architecture search looks for a smaller model that is equivalent to (A) and suitable for devices (B). A team decides a budget based on how much more efficient model (B) must be to deploy. Practitioners use techniques to compress their model (B) to reach the budget (C).
A diagram showing three ML models that shrink over time after ML developers apply (1) architecture search and (2) model compression to finally produce the most efficient model possible.
5.4.2. Budgeting by User’s Experience of Model
Many aspects of budgeting come directly out of product feature design for where, when, and how often a model will be running.
[Uncaptioned image] [Uncaptioned image] One-Time, Real-Time, or All the Time?
Latency budget is almost always a per-feature UX decision dependent on how a user perceives the model output and how often the model needs to run. For example, when tagging people in a new photo, the model needs to work fast enough that users do not notice a delay. Models on-device remove the wait on network response times, which make these latency budgets easier to achieve [P2].
Models that only need to run once generally have a looser resource budget than continuous models. For models running in real-time, only a few milliseconds per evaluation may be the maximum budget [P1]. This is most crucial in real-time scenarios where the user can perceive the model working, which often occurs in live video applications [E13]. A user will notice if their video feed appears choppy or a background filter appears delayed. Latency budget needs to be as low as possible. Equally, if not more, strict latency budgets are given to models that need to run all the time, i.e., “always-on”. Always-on models are continuously using the device’s resources, so they are required to be as quick and lightweight as possible. Some always-on models require a timely response to the user, for example, an always-on model that listens for a user to trigger a voice-assistant.
[Uncaptioned image] User Opt-In or Background Task?
There are some applications where a ML model is the main enabling technology. For example, if the user explicitly asks a voice assistant for help or uses an app to translate languages. If a user explicitly presses a button to trigger a model, the user is in control of when the model runs (and does not run). Product designers will allow models a bigger budget for power and compute resources when the user explicitly opts-in [E11, E7, P2]. It is also a common scenario to have models running support functions in the background. Similar to any other background task, these models require a strict budget to stay unobtrusive:
“If the user will never perceive the model, we still want the memory footprint to be inconspicuous so it doesn’t interfere with them using their device.” — E7
[Uncaptioned image] Can it Wait for a User to Sleep?
Many devices today will download and install their routine updates at nighttime, or whenever the device’s user sleeps and has their device connected to power. Updates are intensive computational tasks that can be disruptive to normal device function, thus it is conventional to wait until a user is not actively using their device. By the same logic, scheduling a model to run overnight is a good option if the model’s output is not something needed immediately (e.g., updating tags of a photo library). If the user’s device is plugged in, this also alleviates the concern of a model draining the battery. Disruption to the user is minimized, so overnight models are also free to take a longer time with a bigger resource budget [E13, P3].
Strategy #1: First estimate the accuracy and latency of the original model architecture you want to run on-device. Then estimate how often the model needs to run (less often is “easier”) and what it needs to deliver (less accurate is “easier”) to budget the worst-case performance that will be acceptable for your user experience. The gap between your starting estimate and worst-case budget will tell you how feasible your on-device ML plan is.
5.4.3. Top-down Budget by Application & Device
As illustrated in Figure 2, an application or device has finite resources—and ML models are only one component of a system. For this reason, major budget allocation is typically decided by people with far reaching views and ownership in an organization [E1, P1, T1, P9]:
[The product lead] looks at budget for whole feature and allocates budget to individual models from there. There are dozens of algorithms. [Each model] gets budget based on priority, practicality, and executive decisions on what is the most important algorithms to give space/time to.” — T1
The difficulty of hitting budget can vary enormously depending on the model. Thus negotiation over budget allocations can continue and evolve until late in development [P1].
[Uncaptioned image] Size Budget
Any one application should not take up too much disk space on a device. The challenge is that as more and more new features make use of ML, there are more models to fit in the same amount of storage:
“Because of [neural networks], and the rising popularity of huge transformer models, compression becomes more and more important.” — E9
Figure 2. Models needs to share space with the rest of an application. Model budget typically reflects how “valuable” a model is compared to all other features of the application.
A rectangular bar representing a fictional application’s storage size that is broken up into various pieces. The piece that represents how much storage ML models take up is highlighted in blue.
[Uncaptioned image] Power Budget
A model’s power budget is a measure of how much it is permitted to drain a device’s battery life. Power consumption of a model can be tricky to estimate, and should be measured empirically on the target device for the most reliable estimate [E6, T4, P9]. Once again, real-time or always-on models require much stricter power budgets, because they continuously draw power. Instead of measuring power at one moment of execution, these continuous models can be thought of in terms of the total percentage or total minutes of battery they consume in a day [E6, P3, E7].
[Uncaptioned image] Thermal Throttling
Closely related to power consumption is device temperature. In the words of E7: “ship high-quality ML models that don’t melt the user’s device.” By melting, E7 refers to both a device that is uncomfortably warm to the touch, and a device that has been forced into thermal mitigation. Thermal mitigation, or thermal throttling, is when a device slows down processing to protect itself from overheating. Hot and slow devices is a terrible user experience that must be avoided. Thus, memory and power budgets must be set according to thermal measurements [E6, E7, E4, E1, E2]: “Heat throttling is everyone’s budget!” [E4].
Multiple Models in Concert
A development team may measure the power, latency, or memory load of a single model as they develop it, but testing a single model in isolation is insufficient in situations where multiple models are powering an app or experience at the same time [E6, E13, E10]. Multiple models running simultaneously will affect the overall memory pressure experienced by device at any moment, and thus the latency of each model as well as total power consumption. For these situations, teams need to experimentally test models running in concert to get accurate measures of total consumption, and experimentally refine their budget for each individual model from there [E13, E10]:
“You could use too much memory, and that’s a no-go. You don’t know that until you test on device. [The memory bound] is not something you can compute ahead. There are multiple models happening at once, so until they are all put together you can’t see exactly what the memory is going to be on a process.” — E13
Strategy #2: Budget a model’s resources by it’s value-add to your overall experience. For a reliable estimate, run your model on the actual target devices early and often during model optimization. Measure compute usage, memory pressure, battery consumption, heat, and your model’s storage size relative to the overall application. Run the model alone and in realistic, high-load scenarios where many process are running simultaneously to find out if your model is too resource greedy. Remember to not be a bad citizen: an ML feature that does not respect the resource needs of the overall device will result in a poor user experience.
5.4.4. Edge, IoT, and Low-power Devices
Devices like laptops, tablets, and mobile phones have far more memory and power resources than smaller, low-power “edge” devices, such as wearables or IoT devices. For edge devices, budget is tight enough that even the bytes of the modeling code text itself can matter [E11]. For these devices, an expert recommended avoiding neural networks all together due to their high cost, and approximate the same accuracy as much as possible using the lowest profile conventional machine learning algorithms, such as decision trees [E11].
5.4.5. Gating & Variable-Budget Options
There may be features where it is not (yet) technically possible to fit the ML component entirely on device. One option is gating, where only some inference is done on-device: A small simplified model lives on-device and is only highly accurate for common inputs. If the small model detects complex or uncertain inputs, it invokes a high-powered ML model that uses more resources, or it sends the harder input to an external model on a server (see background Section 3.6.2) [E11].
Another option is variable budgeting. Product teams may be able to keep essential parts of a feature running smoothly under intermittent network availability by dynamically changing whether it uses higher-accuracy models or lower-accuracy models. Similarly, product teams can also choose to lower ML accuracy when needed to keep a feature running under high-device memory load:
“So people think about model compression as static resources, but the actual resources on device are dynamic based on what’s running. So it’s useful to be able to on-the-fly change how much resources your model is uses.” — E10
For any model, there will be some size and efficiency budget at which the model can be shrunk without any noticeable difference to the user. Below that threshold, accuracy degradation may degrade UX. A budget must balance the best model accuracy with the lowest possible resource footprint.
Strategy #3: On-device ML does not need to be all-or-nothing. If current compression techniques cannot fit your ML feature on-device, try breaking the feature down into subtasks. Delegate smaller subtask models to live on-device and delegate larger ML workloads to a server. Prioritize on-device ML for feature subtasks that will help preserve your user’s privacy and keep critical functionality responsive in the absence of a network connection.
5.5. Considerations for Investing in Compression
Participants strongly emphasized that model compression is an investment. Compression techniques can take intensive engineering effort to apply—and do not always work. There are many circumstances that can cause one compression technique to generate enormous savings in one model, and completely fail to produce any savings in a different model, potentially after weeks of wasted engineering effort. While this work is not an exact science, we detail general tips and guidance from practitioners. Note that most of these techniques assume the model to compress is a neural network.
5.5.1. Maximize Architecture Savings First
As shown in Figure 1, practitioners often start optimizing a model by finding a smaller architecture first. “If you’re running a common ResNet, look at strides, layer widths, number of layers” [T5]. Participants also recommended neural architectures designed for device efficiency [E13, P5, E3, E11, T6], such as MobileNet (Howard et al., 2017). Some applications do not need a heavy neural network; instead it might be possible, and even beneficial, to convert into a simpler decision tree or SVM [P3]. Architecture savings are seen as a more reliable first approach before attempting compression [E13, E1]:
“If you’re trying to reduce the size, model compression is effective, but more like a last resort. If you have a model that is just overkill, [compression] can shave off 20-40%, but just changing architecture to something smaller will have way bigger impact.” — E11
Due to the high training cost of some models, participants discussed manual architecture selection (as opposed to auto-ML solutions that automate various stages of ML development) based on their expertise: “It’s really trial and error by hand monitoring the hardware usage [and] identifying bottlenecks,” [E2]
5.5.2. Check the Architecture Against the Hardware
A critical caveat to architecture optimizations is that it is easy to be fooled. A common mistake is to trust that academic efficiency results will be reproducible: “Just because you read a paper, don’t expect what is “efficient” in a paper to exactly work in practice,” [E7]. The issue is that at the lowest level of computation, different hardware optimizes for different operations. Efficiency measures reported in papers are only reliable for the specific hardware set they were run on. Even for specialized accelerator hardware designed for ML models and neural networks, the hardware will likely only support certain types of operations. Since hardware cannot be altered at the same pace as software, the latest neural layer architectures from papers may not be possible to run efficiently on-device. Practitioners emphasized the need to adapt their neural architecture to better fit the hardware [T5, E7, P9].
5.5.3. Test Against a Range of Hardware
In conventional machine learning work, and more generally much of modern software engineering, it is not required to know the details of the hardware that code will run on. In efficient machine learning, hardware may be one of the greatest challenges when working in on-device machine learning:
“You must consider different layers of abstraction of a model from code to hardware. This can be hard for developers to understand.” — E6
To further complicate matters, many times practitioners are not considering a single specific hardware implementation, but rather a class of hardware. Classes of hardware include platforms such as mobile phones, smartwatches, tablets, computers, and others. Consider a practitioner building a model for a smartphone. There are many types of smartphones, and versions of existing smartphones each with their own memory and compute details. Hardware also changes over time, and practitioners may need to consider older versions of hardware that have already been shipped. In the same way that front-end developers build responsive UIs and applications for multiple screensizes, so too do ML practitioners now need to optimize and test multiple hardware implementations specific to a set of devices [E6].
Strategy #4: If starting from scratch, chose a model architecture specifically designed for mobile devices, but be aware that an architecture may not perform as-advertised due to implementation differences in hardware. It is crucial to profile models on your physical target hardware—and for every hardware version you aim to support. Older devices usually have fewer compute resources for your model.
5.5.4. Accuracy v. Effort Trade-Off
Practitioners we spoke with had a sense of “accuracy v. effort” [E6] or “risk v. reward” [E13] for some of the most common compression approaches. As a general heuristic, compression techniques that need little-to-no retraining will be cheaper to apply—but at the cost of steeper model accuracy degradation. Since some of these models take days to train, involving complex compression logic in the training process is costly. Practitioners must consider cost in money, time, and effort that implementing a specific compression technique could take.
(\$) Post-training Quantization
Quantizing a model’s weights after the model has been trained was widely considered the first go-to compression technique for many applications.
“So quantization is a big one. You can usually quantize to 8-bit integers without losing accuracy. This isn’t always the case. You do need to be careful where you apply quantization. But generally it’s a pretty generic technique across different hardware. You can usually cut from fp32 to fp16 and can get speed up by going to integers.” — E4
Since weight quantization can be done without additional training, it is considered cheap. 10/30 participants discussed using this approach [E6, P8, E9, T3, T5, T1, E13, E7, T2, T2]. Although post-training quantization is considered “easy” [E9] as far as ML compression techniques go, practitioners emphasized that it still often takes complex code to implement and there are many algorithm variations (Gholami et al., 2021) to experiment with [T5]. For models that need high accuracy, post-training quantization may not be enough to hit budget without unacceptable accuracy degradation [E9, E4, E13, E5].
(\$\$) Compression with Fine Tuning
When model accuracy goes down after quantization or pruning, fine tuning can help recover accuracy (Gholami et al., 2021; Hoefler et al., 2021). Fine tuning a model costs training time, but is necessary in many situations to recover accuracy. Pruning techniques almost always require fine-tuning on the pruned model (Hoefler et al., 2021). 6/30 participants discussed successfully utilizing this approach [P8, T3, T1, E8, E13, E10].
(\$\$\$+) Compression-aware Training
Though it introduces additional complexity to model training, compression can be approached as yet another optimization that a model needs to learn during training. 11/30 participants discussed having experience with quantization learned during training [E6, E9, T5, E3, E5, E7, E10, E1, E12, T2, P10]. This approach is generally considered best-in-class and often required when designing always-on ML experiences that require low latency, e.g., real-time computational photography models. Practitioners advised training-aware compression when the model needs to be compressed significantly to meet budget while keeping high accuracy:
“If you want to go to lower bit quantization, such as 4 or below, it’s almost impossible to use post-training quantization because the difference in accuracy gets way too big. So for this level of compression you need to do training-aware compression.” — E9
Practitioners had less experience with learned sparsity techniques, calling learned sparsity “high risk high reward” [E13] because it is much harder to control than post-training sparsity techniques. Although training-aware compression is considered the best form of compression for the greatest model shrinkage with the least accuracy degradation (Gholami et al., 2021), a major drawback is that is must be planned for with initial model training: “Not starting early with compression is a dead end,” [E3].
Practitioners use their experience and early testing to judge how much compression a model will need. For example, large vision models may start far exceeding their size and power budgets, and require heavy compression to meet budget. In these scenarios, training-aware compression is likely necessary, which requires planning ahead: “It has to be done from day 0, not day 100,” [E7].
For other applications, practitioners suggest estimating how much compression will be feasible with simple post-training quantization. To estimate quantization savings before training a model, first initialize the ML model architecture with random weights, then quantize, and test the model’s speed and size on-device. Even a coarse estimate may be worth getting a sense of the magnitude of savings from quantization [E13]:
“[It] gives you a sense of how much quantization is going to help at all before you go through expensive training step.” — T5
If post-training techniques like quantization produce results that are too far from budget goals, teams can then pivot to training-aware approaches.
Strategy #5: Before you heavily invest in a particular compression technique, start with a cheap estimate of how much savings you might expect to gain. For example, initialize a mobile-friendly architecture with random weights, then quantize it. Profile this “stand-in” compressed model on-device and compare its efficiency to your budget. If it meets budget, then quantization may be enough. If it is far off budget (such is the case with many real-time computer vision models), consider more intensive training-aware compression techniques.
5.5.5. Determine if a Model is Memory Bound or Compute Bound
Another useful strategy is to consider ahead which metrics are likely to be the biggest issue for a model type [T5]. A model is memory bound if the size of the model or the size of the data is the biggest issue. A model is compute bound if the speed/latency of the model is the biggest issue. For instance, “Computer vision tends to be memory bound and not compute bound. It’s fast but the data sizes are massive” [T5]. Participants also suggest going layer-by-layer to identify bottlenecks since individual layers can be either memory or compute bound individually [E2].
If memory-bound, architecture changes, sparsity techniques, or Palettization techniques can help reduce size on-device. If the data size is causing the memory issue, reducing data resolution for model processing can help [E8], e.g., reducing video data from 1080p to 720p. On-disk storage is considered the easiest memory savings to get from model compression [E13].
If compute-bound, techniques like quantization can help speed up and simplify complex matrix math. Quantization can also save memory since smaller numbers help keep computation in cache memory [E12]. Cache locality is a common latency bottleneck to check [E6, E12], that also helps with power problems:
“Latency is easier than power to improve. You can check cache locality, moving data takes a lot of power so that’s a big one to check.” — [E6]
5.6. Evaluating Compressed Models: Efficiency, Accuracy, and Artifacts
The goal of model compression is to reduce a model’s size while preserving its accuracy, or what some refer to as “acceptable accuracy degradation” [E8]. In general, past a certain point, shrinking a model will likely degrade its accuracy. In literature, there are examples showing that compression can actually improve accuracy, by acting as a model regularizer and forcing the model to generalize better (Kusupati et al., 2020; Touvron et al., 2021). This phenomenon is sometimes referred to as Occam’s Hill (Hoefler et al., 2021): as light compression is applied, there can be an slight accuracy bump as the model is forced to generalize; however, as one increases the compression strength the model becomes too general to properly work and accuracy quickly drops. While these results have been observed on academic benchmarks, practitioners said that their models tend to generalize quite well already, so most often they see little improvement from the regularizer effect of compression [E8, E1]. In addition, many practitioners start with very large models they need to shrink models past the point of observing accuracy boosts from compression. Thus, we focus this discussion on the much more common scenario where practitioners must wrestle with accuracy degradation.
5.6.1. The Trade-off Curve Visualization
To empirically compare multiple compressed models and select the one that satisfies budget requirements, practitioners typically plot a model behavioral metric (e.g., accuracy) on the y-axis against any performance metric (e.g., model size, model latency, or power consumption) on the x-axis (Vasu et al., 2023b, a). This is illustrated in Figure 3. Each dot represents a model architecture and compression pair. These charts are called multiple names throughout different teams, including the trade-off curve, the an accuracy Pareto curve, or the tuning curve.
P3 described a scenario when optimizing over 6,000 small models. In their application, the ML model was already on the order of kilobytes; therefore, they could retrain many new models quickly. Comparing thousands of models in parallel, they plotted the F1-score on the y-axis against the model size on the x-axis. This arrangement formed the usual curve. Teams then filtered out models that do not satisfy their budgets, and ultimately selected a single models in the “knee” of the curve that has the best balance between what metrics they care about (e.g., Pareto optimal).
Figure 3. A common chart used for model comparison and selection, where the y-axis is a model behavioral metric (e.g., accuracy) and the x-axis is any number of performance metrics (e.g., latency or power consumption). Each blue dot represents one model trained with different architectures or compression schemes, and the red line indicates the accuracy budget threshold. Practitioners look for the “knee in the curve” across charts: choosing a model above the accuracy threshold that minimizes latency and power consumption. In these charts, the selected model is colored red.
Two scatter plots where each dot represents a compressed ML model. The y-axis encodes model accuracy from 0%–100%. The x-axes of each chart encode latency (in ms) and power consumption (in mw) respectively. Each chart shows a slight curve from the bottom-left to the top-right, showing a trend that as a model gets more accurate it tends to have a higher latency and larger power consumption. The models that are chosen to be the best are the in knee of the curve and balance accuracy, latency, and power.
5.6.2. Compounding Changes Degrade Accuracy
A common challenge with maintaining accuracy is that model compression techniques can compound in unintended ways. In some scenarios, small optimizations throughout a model build on one another, so by the time a data input reaches the end of a network, its prediction is totally off. In ML this phenomenon is similar to the concept of exploding/vanishing gradients (Bengio et al., 1994; Pascanu et al., 2013). The effect of this compounding error is not obvious to spot in beginning [E9].
“When you have addition that takes two quantized values to a quantized output, it’s not easy to check that the range of the inputs are the same as the output ranges. You tend to lose resolution in the output where one branch dominates the range, and you lose the range of the lesser branch. You need additional care to check that they have the same range tracking during the forward quantization pass.” — E6
For quantization, practitioners advise a “gut check” to ensure the quantized weights and activations match the ranges of the original model, or else accuracy will be badly degraded [E6].
5.6.3. Robust, End-to-end Data Evaluation
Since the amount of success in model optimization varies significantly by architecture and task, “without a clear evaluation strategy you won’t know if you’re making things better or worse,” [E11]. Some ML-powered experiences are composed of multiple models working together. Participants said they first test models individually to ensure they work standalone, then evaluate the multi-model systems to test compound effects [E10]. Practitioners said the curve visualizations (Figure 3) are often the output of evaluation pipelines.
Many teams do extensive model evaluation, error, and failure analysis [T2, E12, P11, P6, P7, E10, T4]. “We cannot assume compression doesn’t change model behavior, so we look at confusion matrices and instances where model gets it wrong” [E12]. In one project that P11, P6, and P7 worked on together where small mistakes could lead to a poor user experience, they built “unit tests” for different curated testing datasets to monitor their accuracy over time. E6 also emphasized using unit tests of data to observe how single instances move through the model to both understand how much computational change occurs after compression and make the necessary adjustments to the model code. “Data so influential to hitting accuracy targets” [E3].
Strategy #6: Compression can degrade the accuracy of a model and change its behavior in unpredictable ways. It is essential to create a robust evaluation pipeline (e.g., defining metrics, curating test sets) before you start optimizing your model, so that you can reliably observe shifts in model error afterwards. To prevent degradation from a failed optimization, compare optimized models with varying amounts of compression to your original model, inspecting the metrics, subpopulation behaviors, and internals, such as weights and activations, to ensure they are within expected ranges.
5.6.4. Model Compression Artifacts
As with other types of media, such as image, audio, and video data, if too much compression is applied it can produce compression artifacts: noticeable distortions in the media (Figure 4). For example, compressing an image too much can produce blurry and blocky shapes over the subject matter; compressing audio can change the sound quality and waveform of the music. Multiple participants describe scenarios where their model had artifacts, although they did not borrow this language from other types of media:
“Compressing like 90% can make things unstable with strange side effects. It’s hard to figure out when you compress too much.” — E13
It is tempting to think of accuracy degradation as the only ML artifact; however, there are more subtle, even sinister implications depending on what a model is used for. For example, where some subpopulation of data is underrepresented in a dataset, e.g., at the tail of a distribution, it could be that ML compression techniques removes this tail, amplifying existing biases. Examples of this have been observed empirically in the few publications investigating robustness and evaluation of compressed models (Hooker et al., 2020, 2019; Ogueji et al., 2022; Liebenwein et al., 2021; Gale et al., 2019).
Figure 4. Illustrative examples of compression artifacts in images and audio, but what do machine learning model artifacts look like and how do we identify them?
A grid of images showing the effect of too much compression which creates artifacts on images, audio, and ML models. The image artifacts show a reduction in colors and distort the image. The audio artifacts show missing frequencies from the audio’s frequency range. The ML model is a small neural network with a question mark, indicating that we do not know the full scope of what ML compression artifacts look like.
One specific compression artifact example story was told from multiple participants. In the case of an object detection model that needed to run at a high frame rate, during development teams noticed one day that the bounding boxes were jittering in their demo. This finding was surprising, since the metrics from their most recent model iteration were reporting no changes, but “there were some weird side-effects” [T5]. It was not until someone debugging the problem realized that they had applied quantization throughout the neural network, including the final prediction layer. This coarsening of the data at the output produced correct bounding boxes, but resulted in a poor user experience. Another participant had seen this before, and remarked if the output of a model is “fine grain, don’t quantize” [T1].
Generalizing from this example, one learning that teams have found is that when a model’s prediction needs to be continuous, smooth, or user-perceivable at a high frame rate, it is best not to compress the final output layer (Deng et al., 2020). Participants told us that other modeling tasks can to be harder to optimize than classification [E5]. “Regression models have been an absolute nightmare to compress” [E7]. Anytime the output needs to be a continuous or floating point value, then compression techniques like quantization can produce artifacts [E12].
5.6.5. Demoing User Experiences
ML-powered experiences can be dependent upon multiple models working in concert, and the “best way to test compound [model] effects is having end-to-end evaluation” [E10]. Another strategy to catch artifacts is testing models “as close to the metal as possible,” i.e., loading the models on-device and building demo applications to run outside of lab environments. “Since different hardware are not bit-wise accurate, metrics won’t capture these changes” [E12]. Another practitioner told us “user experience differences are the really important cases to isolate” [P5]. To find these user experience changes, teams have prioritized building demo applications where different models, e.g., a baseline model and a compressed model, can be dynamically toggled back and forth for testing [T4]. Some of these demos toggle between a model running on-device and a model running on a server [E12]. Interactive and live demos like these allow for ML teams to get feedback from other product stakeholders [E10]. They are particularly helpful for designers to get an overall “feel” of a model: interacting with a model in its intended environment to understand its capabilities and limitations [T4]:
“Get [the experience] in hands of people as fast as possible.” — T4
This notion of the “feel” of a model was described multiple times. If an “ML engineer retrains [the model] and says it’s better, we still need see if it feels better,” [T4]. To try and attribute the feel of a model to actionable development steps, practitioners show debugging modes on-device to observe live, real-time charts of a model’s prediction and confidence, to help find edge cases and drill into specific errors that metrics may or may not capture [P11, P6, P7].
Strategy #7: Despite all the effort to create criteria and metrics to quantitatively measure and benchmark your model, your evaluation pipeline may not capture every aspect of your model’s performance and behavior. Even if your model passes evaluation, move it on-device, in the intended environment in which it will run, and get it in the hands of users to demo. There is no other way to capture the feel of an ML-powered user experience.
6. Design Opportunities for Efficient Machine Learning Tools and Interfaces
Given the unique challenges to efficient on-device ML, where can human-centered ML researchers, practitioners, and designers start to engage? As efficient machine learning techniques are driven forward by advances in hardware engineering and ML research, there remains a major barrier in practically applying these techniques for real-world ML features and user experiences. Tools influence what is possible: “there is a gap between what is possible with machine learning and what [tools] are being used,” [E4]. Currently, the tooling for efficient on-device ML is underdeveloped, “homegrown and ad-hoc,” [E9]. Moreover, current tools focus on individual algorithms (subsection 2.6) instead of the holistic process: “this a newer area, many tools are specific to a project and tend to be prototypes to prove feasibility,” [E9]. As we have demonstrated in this paper, there is considerable design planning, experimentation, and strategy that experts use to make on-device ML a reality. Proper tools could help guide and educate practitioners to develop these skills:
“Tooling is education products disguised as software. Better tools make it easy to do correct things and harder to do incorrect things.” — T2
We next share interdisciplinary HCI + ML research directions for supporting practitioners. While the space of tooling opportunities is wide and will evolve over time, we provide a few actionable recommendations for future tools and interfaces that are ready for development today.
6.1. Developing Intuition for Compression
Although modern libraries for machine learning make it easier than ever develop models, that does not mean a practitioner know how to best train, evaluate, and deploy models. Machine learning is an inherently iterative and and empirical practice (Patel et al., 2008), and developing intuition for how models learn and behave is a major competitive advantage over blindly tweaking hyper-parameters. Interactive playgrounds that provide a safe environment where practitioners can quickly build and test their ideas could be a great onboarding experience for learning about efficient ML and model compression. Existing examples include TensorFlow Playground (Smilkov et al., 2016) for small neural networks, CNN Explainer (Wang et al., 2021c) for convolutional neural networks, and GANLab (Kahng et al., 2018) for generative additive models. An efficient ML playground could allow practitioners to learn about and build intuition about different compression techniques [E12], model architectures [E7], and their effect on hardware combinations [P9]. For example, comparing two compression techniques and inspecting their impact on metrics such as power, latency, and model size.
6.2. Comparing Models Across Compression Schemes
A common scenario in efficient model development is considering the trade-off between accuracy and performance, as discussed in Section 5.6. While conventional ML development requires model comparison between architectures and hyper-parameters, the additional optimization metrics that practitioners must navigate add additional complexity in efficient machine learning [T5]. Both accuracy and performance encompass numerous metrics to balance between. Better tooling to support model comparison could help practitioners compare a feasibility, or baseline model, against different compressed models to select the best model possible.
6.3. Finding Model Bottlenecks for Targeted Compression
In Section 5.5.5 and Section 5.6.4, practitioners discuss analyzing a model layer-by-layer. Targeted compression is the practice of selectively optimizing certain components of a model, for example, larger layers in a neural network. Practitioners described the operations that go into targeted compression. Currently the process is tedious, manual, and time-consuming, where “you must go layer by layer, operation by operation,” [P11, P6, P7]. There exists a space of interfaces that could help practitioners look inside their models, by mapping metrics to specific layers of a network [T5], finding the performance bottlenecks in a network [E5], comparing the input and output of these bottleneck layers [E9], and selectively optimizing them.
Targeted compression tooling may also ease the difficulty of thinking on the hardware level. For example, once an on-device model that is written in Python is compiled onto specialized hardware, many of the operations look far more low-level than Python code. To complicate matters, low-level hardware operations may not have a one-to-one mapping back to the original Python code, due to optimizations in the compilation process. Better tools could help practitioners perform analytics on their models to find bottlenecks while helping navigate between different layers of abstraction.
6.4. Evaluating Multi-model ML Systems
It is often assumed that model evaluation is done on a single model, isolated in a “lab environment.” In practice, this is not the case. Not only are models integrated into a larger app or codebase, many modern ML-powered experiences are the result of multiple models working together. This makes evaluation when models are arranged in chains, where the outputs of one model are the inputs of another [E10]. If you compress one model, how does that impact downstream models? Recalling the concept of compounding error discussed in Section 5.6, small errors introduced in earlier models can compound to produce bigger errors by the end of a model chain [E10]. E3 emphasized this challenge, saying that the future of ML driven user experiences will be accomplished through multi-model systems. When multiple models are running simultaneously, each can be evaluated individually, but also must be considered as a whole, which makes it “super hard to reason about,” [E3]. Future tools that generalize and can evaluate multi-model machine learning systems will have a big impact in helping practitioners build, debug, and make sense of large data-driven applications.
6.5. Easier Testing on Hardware Environments
One recommendation repeated again and again by practitioners in this study was to measure model performance on device. The performance of one model will like differ across different hardware [T4]. However, testing on real hardware can be a major barrier. Due the variety of mobile devices and different versions of hardware, it can be hard to build for multiple hardware implementations simultaneously. The average ML developer (outside of a hardware company) may not have access to all versions of all devices their users might have. When it comes to future tooling opportunities, E6 says it best, “the tools need to expose and take care of the cases for different hardware to maximize the hardware efficiency.”
6.6. Automating (Some) Compression Experiments
While many of the practitioners we interviewed discussed creating efficient models as an human-in-the-loop iteration process, there are opportunities to automate applying different compression schemes and present results to developers. An apt analogy is what AutoML is to hyperparameter searching: instead of sequentially trying different compression schemes, perform a grid search and parallelize many different tests simultaneously. P3 is excited by tools that could automate certain parts of a model, perform some analysis to figure out which techniques work and which do not, and finally present a summary of experiments make suggestions for the best fit compression schemes. However, we do not expect the process of creating efficient models to be completely automated. Throughout our study, experts made it clear that successful model compression is an iterative development process, but leveraging the strengths of automation where appropriate suggests rich mixed-initiative approaches to creating efficient models.
7. Limitations & Validity
Optimizing and creating efficient models to run on mobile devices is still relatively new, and best practices are evolving alongside the rapid pace of ML research. Thus we expect not all the practices mentioned in this paper will stand the test of time. As discussed in Section 4, studying experts from a single organization naturally limits the perspective of this paper, therefore we also expect certain details of our findings may not generalize. Nonetheless, we believe that the high-level strategies for on-device ML shared in this work—including budget design, strategies for investing in model optimization, and proper compression evaluation—will generalize to be useful for many different kinds of devices, domains, and ML user experiences. We are eager to watch how this interdisciplinary area of research progress in the future and reflect back on the practices outlined in this work to see which have remained, been removed, or reinvented.
8. Conclusion
In this research we illustrated some of the pragmatic design considerations that go into efficient on-device machine learning. While this paper is far less technical than related work from the ML literature, we feel the interdisciplinary bridge to UX and product design aspects are critical to bringing on-device ML into more approachable and popular practice. Through the results of this study, we are able to spotlight the holistic, end-to-end considerations that world-experts in model compression have developed from bridging ML research into real-life deployments of on-device ML experiences.
Acknowledgements.
We thank our participants at Apple for generously sharing their time and knowledge. We also thank the many colleagues that have given their feedback on the paper. Specifically, we thank Kayur Patel who sparked early ideas for this line of research, as well as Mohammad Rastegari, Sachin Mehta, and Yannick Assogba for their advisement on this work.
Appendix A Interview Questions
The list of questions prepared for each interview participant.
Background ML Information
  1. Q1.
    What is your team?
  2. Q2.
    What is your role?
  3. Q3.
    How many years of ML experience do you have?
  4. Q4.
    How many years of efficient ML experience do you have?
General Compression Questions
  1. Q5.
    Describe the overall use case for model compression in your work. What are your main motivations to use model compression? Why do you care about model compression?
  2. Q6.
    Model compression details:
    1. Q6.1
      Which model compression techniques do you use, and why? Are there trade offs or preferences?
    2. Q6.2
      Do you do compression as a final step or as a part of the training process?
    3. Q6.3
      In your experience, are there any pitfalls people applying compression should be aware of?
    4. Q6.4
      How did you figure out your budgets and how did you satisfy them?
  3. Q7.
    Do you compare compressed models against baseline models? If so, how, and what do you look for, e.g., power and performance, other metrics, or user experience changes?
Compression Tooling Questions
  1. Q8.
    What tools or visualizations do you use in your compression work? Please be specific, e.g., specific charts, views, or metrics. How do you use these tools?
  2. Q9.
    What do you like about these tools?
  3. Q10.
    What features or future tools would help you conduct better model compression work?